Drives - n-More Manufacturing manufactures 256GB SD cards (memory cards for mobile phones, digital cameras, and other devicesRequirement 2. What would the companys monthly operating income be if the company sold 140,000 units? Use the following tablRequirement 4. What is the breakeven point in units? In sales dollars? Begin by identifying the formula = Breakeven sales inRequirement 6. Management is currently in contract negotiations with the labor union. If the negotiations fail, direct laborRequirement 9. What is the companys current margin of safety in sales dollars? What is its margin of safety as a percentageGiven this sales mix, how many of each type of SD card will the company need to sell to reach its target monthly profit of $2

positive rating given upon completion thank you!
Answer #1

1.Contribution Margin per unit = Selling price per unit – variable costs per unit

= 25-7.6-6-4.40-3

= $4 per unit

CM % = CM/Sales

= 4/25

= 16%

Total CM = 4*110,000 = $440,000

2.Operating Income = Contribution Margin – Fixed costs

= 4*140,000 – 279000

= $281000

3.Operating Income = 4,500,000*16% - 279000

= $441,000

4.Break even point in units = Fixed costs/CM per unit

= 279000/4 = 69,750 units

In Sales Dollars = Fixed costs/CM ratio

= 279000/16% = $1,743,750

5.Units required = (260,100+279000)/4

= 134,775 units

6.Break even units = (279000+22,500)/3.4

= 88,676.47 units

7.Operating leverage factor = CM/Operating Income

= 440,000/161000

= 2.7329 times

8.Operating income will increase by 6%*2.7329

= 16.3974%

9.Margin of safety = Sales – break even sales

= 25*110,000 – 1,743,750

= $1,006,250

% = 1,006,250/2,750,000 = 36.59%

10.Weighted average CM

256 GB

512 GB

Total

Sales price per unit

25

50

Less: Variable cost per unit

21

20

Contribution margin per unit

4

30

Sales mix

2

1

3

Contribution margin

8

30

38

Weighted average CM

12.6667

Break even units = 279000/12.6667 = 22,026.31 units

256 GB = 14684.21 units

512 GB = 7342.10 units

Units required of 256 GB = [(279000+260100)/12.6667]*2/3 = 28,373.68 cards

512 GB = 14186.84 cards

Lower, since weighted average contribution margin has increased due to 512 GB Cards